# Exploiting the diversity of modeling methods to probe systematic biases in strong lensing analyses

### v1

@@ 8.2 Model degeneracies between the lens and the source
As presented in Sect. 6.4, we observed correlations between the level of recovery (the relative error) of lens and source properties. In particular, we find a correlation between the lens power-law slope $\gamma$ and the source effective $r_{\rm eff}$, with correlation coefficient $r=0.9\pm 0.4$, namely a positive correlation with a statistical significance of $2.3\sigma$. It is well-known that correlations between the source scale and the mass density profile of the lens can arise, which are often seen as a manifestation of the MSD (Falco et al. 1985). The MSD originates from the mass sheet transformation (MST), and its consequence is as follows: the addition or subtraction of a infinitely thin mass sheet from a power-law mass density profile changes, to first order, the slope of the density profile at the Einstein radius (e.g., Schneider & Sluse 2013; Blum et al. 2020; Birrer et al. 2020), while rescaling the source proportionally. A positive mass sheet locally increases the density slope, which in turns induces an increase of the source size through the MSD. This effect has also been empirically explored in Birrer et al. (2016) by using the explicit scale encoded in a source model based on shapelets. The positive correlation we find between the biases in $\gamma$ and $r_{\rm eff}$ may thus be the signature of the MSD: all the lens models we consider infer slightly too large density slopes and source effective radii.
@@
There is a general trend of correlations between source light shape and lens mass shape that we observe in our results. For instance, we note correlations between the recovery of $q_{\rm s}$ and $q_{\rm m}$ or $\phi_{\rm ext}$, $r_{\rm eff}$, and anti-correlations with $\phi$ and the lens centroid. We also observe strong correlations between the error on $q_{\rm s}$ and that of the lens centroid. These correlations are more challenging to unambiguously interpret compared to those associated to the MST. We argue that they may be related to the source position transformation (SPT) outlined in Schneider & Sluse (2014) and further developed in Unruh et al. (2017). The latter work effectively shows that transformation of the source profile can be compensated by changes in ellipticity and radial profile of the lens mass distribution. The biases that we observe in our models may be manifestations of this transformation and qualitatively match the expectation for a SPT. Because the SPT leads to an approximate degeneracy (unlike the MSD), it is expected to be broken with higher quality data. It may therefore be interesting to see how these correlations change with higher S/N data or even noise-free mock data.
@@ 8.3 Combining methods to mitigate systematic biases
We quantify the scatter associated with the choice of modeling method in Table 2, which records the mean and scatter of relevant source light and lens mass quantities among all the models. Although dependent on the type of data considered in this work, these numbers provide an estimation of systematic errors that one can expect from modeling imaging data with different methods. The systematic errors we quote here can be interpreted as lower bounds to those of a real-case scenario, since we placed ourselves in an idealized setting—no contamination by the lens light and perfect knowledge of noise properties and mass model family—which removes a subset of the known sources of biases or degeneracies. Ideally, the addition of such complicating factors may broaden the posterior distributions from individual models, thus making them statistically compatible (underlying systematic biases would then be unnoticeable). However, more realistically, the scatter between models is likely to increase as a consequence of more complex models potentially subject to different sources of biases. Performing similar analyses as the one presented here to a wider variety of strong lensing data (different resolution, S/N, lensing configuration, nature of the source, etc.) will help understand the generalization of our results. For this purpose, the framework we have developed should help and encourage the multiplication of such analyses on both simulated and real data sets. However, we show that combining independent methods offers a way to overcome potential systematic errors.
@@
As we show in Fig. 6 and Table 2, combining together the results from an ensemble of methods removes the observed biases. While individual methods display systematic errors in different parameters—and not necessarily always for the same parameters and in the same directions—it is reassuring that overall, we do not observe a significant residual offset after combination. Only for the mass density slope ($\gamma$) one can observe a global trend towards larger values. We quantify in Table 2 the bias reduction between the biases from individual models and the one from the combined model. Among the eight mass model parameters, we find that an average bias reduction of 5.4, which is a substantial improvement on inferences from individual models alone. This result is reassuring and shows that analyzing a given data set using independent modeling methods is an efficient way to mitigate systematic errors.
@@
Beyond the accuracy improvement of a combined posterior distribution over the lens model parameters, the comparison of multiple models—and if possible, truly independent models—is extremely valuable for detecting unknown sources of systematics, updating our current modeling assumptions and techniques and testing additional models with different levels of complexity. Similarly, in source plane, comparing the morphology of different versions of the unlensed object (which is never directly observable) is necessarily valuable for straightening the confidence in a given feature, which in turn improves the robustness of its interpretation. Moreover, different models may predict different observables that may otherwise be overlooked, enabling the potential detection of anomalous lens systems (such as missing images Ertl et al. 2024) or anticipate the detection of additional images unobserved in current observations (e.g. similar to the geometric confirmation of multiple images in cluster-scale lenses, Diego et al. 2023).
@@ 8.4 Extrapolating to time-delay applications
Figure 12: Fermat potential differences evaluated at image positions of a multiply point source located at the the center of the source galaxy. Given that the modeled data (Fig. 1) does not contain the multiply imaged point source, this can be seen as an intermediate scenario between that of a lensed quasar (i.e., centered on its host) and a lensed SN (i.e., has faded away).
@@
We show in Fig. 12 the Fermat potential differences for three hypothetical pairs of lensed images. While the simulated data we analyze in this work (Fig. 1) mimics the case of a lensed SN that faded away, we assume for simplicity that its location in source plane coincides with its host galaxy (although one could also proceed similarly for any other source position that leads to at least two lensed images). We label the lensed images ABCD, order them by ascending Fermat potential (i.e., $\Phi(\boldsymbol{\theta}_{A})$ is the lowest) and consider the three independent pairs $ij\in\{{\rm AB,AC,AD}\}$. In addition to the posterior mass parameters uncertainties, we add in quadrature an additional uncertainty to mimic the limited astrometric precision (for details, see Birrer & Treu 2019), assuming a conservative precision of 10 mas in image plane. The gray shaded area in each panels of Fig. 12 around the true values isolates the typical contribution of the astrometric scatter term, which is in our case small compared to the uncertainties modeling the extended lensed source. As expected from the constraints on mass profile parameters (Fig. 6), we observe a similar scatter around the true values among the models.
Interestingly, we do not observe a clear correlation between model biases on individual lens potential parameters and their resulting biases on Fermat potential differences. For example, the Sparsity+Wavelets model, which display slight biases on some source properties and lens potential parameters (e.g., the lens position), give Fermat potential posteriors well-centered on the expected values. While a robust generalization of these results require further work (e.g., on a larger sample of simulations), the general trend we previously found still remains: the combination of individual posteriors assuming equal weights results in posteriors that are free of biases. These combined distributions are shown with black dotted-dashed lines in Fig. 12, and encapsulate well the Fermat potential difference values. This further gives motivation for time delay cosmography analyses to compare and combine lens models together, in particular those obtained using independent modeling methods (see e.g., Shajib et al. 2022; Kelly et al. 2023). In a follow-up paper, we will explore the dependency of systematic lens modeling errors on the source plane position of the time-delay background object.
@@ 8.5 Application to real data sets
For the purpose of demonstration, we focus on simulated imaging data, but the framework and ideas we present is intended to be applied to real data. The COOLEST standard, which provides a common ground to express lens models from different origins, already supports real data and their corresponding models (see e.g., Galan et al. 2024). Moreover, extending the standard to other types of models (for both the lens and the source) is straightforward. The main complication for applying this framework to real data arises from the ability of applying multiple codes and methods to a given data set. This necessary step obviously requires some level of human expertise and time, as well as computational resources, which can be significant depending on the methods (see also Sect. B).
@@
Although COOLEST significantly reduces the burden to express a lens model in a format that can be readily compared to other models, the acquisition of combined posterior distributions over lens model parameters still requires that at least two distinct lens models are in hand. Here we model the same data using four software packages and six modeling methods, which is likely unrealistic in most real scenarios, especially given the large amount of lenses that are still to be modeled (e.g. in the archival HST data, see Garvin et al. 2022) and that will be discovered in the near future (e.g., Collett 2015). Nevertheless, it is reasonable to assume that it is feasible to apply two or three methods to the same data (not necessarily within the same analysis), since the amount of lens modeling experts naturally grow over time and inference methods are becoming less and less time consuming. Deep learning methods, that we do not explore in the present work, offer promising avenues to accelerate the overall procedure either by proposing preliminary models to be refined with classical methods, or providing additional posterior distributions for final combination, at a negligible cost (ignoring the training phases).
@@ 9 Conclusion
We have conducted a fully blind modeling experiment on strong lensing data simulated with a dedicated software package by one author, while four other authors used four independent software packages to model it. In total, six modeling methods—which differ in their source reconstruction techniques and inference pipelines—have been applied on that same data. We have made a series of simplifying assumptions to keep this novel kind of analysis tractable, in particular regarding the lens light, the lens mass model parametrization and the noise properties. In contrast, the true PSF, lens parameters and source morphology were hidden and the optimization and inference strategies left free to the modelers. We used the lensing standard COOLEST to overcome the challenges that arise when comparing results obtained with different modeling codes. The resulting image and source plane models, as well as model residuals are given in Fig. 2, and constraints on lens potential parameters are shown in Fig. 6. Below we summarize our main results:
@@
  • We investigated how certain model assumptions affect the recovery of lens potential parameters. In particular, we explore (1) how the cuspy nature of the lensed galaxy can lead to systematic errors if the source model is not flexible enough to capture large intensity variations, and (2) how the accuracy of the PSF (the true PSF being unavailable to the modelers) plays a role even for extended source modeling. We find that both an accurately sampled PSF and a source model with large dynamic range (e.g. using a luminosity-weighted source prior) are warranted to reconstruct cuspy lensed sources while minimizing systematic errors on lens potential parameters.
Over the past years, numerous lens modeling methods have been proposed and implemented in different software packages. Here, we selected a subset of those with the goal of using them together, instead of only opposing them. Typically, we refrain from explicitly ranking the modeling methods, which would only be meaningful over an extremely large sample of strong lenses with different data quality and modeling assumptions to ensure proper generalization. In a real-case scenario, we do not have access to the truth, therefore applying and combining different methods together is a pragmatic and efficient way to detect and mitigate systematic errors. As shown, for example, in Suyu et al. (2006) by testing three types of source regularizations, there exists an inherent dependence on the data properties, the lens configuration and the (unobservable) intrinsic source morphology, such that it is unlikely that a single source reconstruction method gives unbiased results in all cases. Our work strengthens this idea and goes further by combining a large collection of models, while investigating their specific effects on inferred parameters. Moreover, we purposefully allowed some level of freedom for the modelers (e.g., masks, PSF, posterior marginalization, etc.), such that our work also illustrates the role of specific modelers’ choices. These choices play a role in the observed scatter between models, and can be marginalized over by combining the methods together.

### v2

@@ 8.2 Model degeneracies between the lens and the source
As is presented in Sect. 6.4, we observed correlations between the level of recovery (the relative error) of lens and source properties. In particular, we find a correlation between the lens power-law slope $\gamma$ and the source effective $r_{\rm eff}$, with correlation coefficient $r=0.9\pm 0.4$, namely a positive correlation with a statistical significance of $2.3\sigma$. It is well-known that correlations between the source scale and the mass density profile of the lens can arise, which are often seen as a manifestation of the MSD (Falco et al. 1985). The MSD originates from the mass sheet transformation (MST), and its consequence is as follows: the addition or subtraction of a infinitely thin mass sheet from a power-law mass density profile changes, to first order, the slope of the density profile at the Einstein radius (e.g., Schneider & Sluse 2013; Blum et al. 2020; Birrer et al. 2020), while rescaling the source proportionally. A positive mass sheet locally increases the density slope, which in turns induces an increase in the source size through the MSD. This effect has also been empirically explored in Birrer et al. (2016) by using the explicit scale encoded in a source model based on shapelets. The positive correlation we find between the biases in $\gamma$ and $r_{\rm eff}$ may thus be the signature of the MSD: all the lens models we consider infer slightly too large density slopes and source effective radii.
@@
There is a general trend of correlations between source light shape and lens mass shape that we observe in our results. For instance, we note correlations between the recovery of $q_{\rm s}$ and $q_{\rm m}$ or $\phi_{\rm ext}$, $r_{\rm eff}$, and anti-correlations with $\phi$ and the lens centroid. We also observe strong correlations between the error on $q_{\rm s}$ and that of the lens centroid. These correlations are more challenging to unambiguously interpret compared to those associated with the MST. We argue that they may be related to the source position transformation (SPT) outlined in Schneider & Sluse (2014) and further developed in Unruh et al. (2017). The latter work effectively shows that transformation of the source profile can be compensated by changes in ellipticity and radial profile of the lens mass distribution. The biases that we observe in our models may be manifestations of this transformation and qualitatively match the expectation for a SPT. Because the SPT leads to an approximate degeneracy (unlike the MSD), it is expected to be broken with higher quality data. It may therefore be interesting to see how these correlations change with higher S/N data or even noise-free mock data.
More generally, approaches that are independent of a specific choice of mass model offer a complementary route to further study the interplay between lens and source structure. As reviewed in Wagner (2019, and references therein), instead of the direct modeling of imaging data pixels, a careful extraction of specific lensing observables can provide constraints onto local lens properties, without assuming global functional form of the lens potential. While such an approach is scale-independent by construction—it does not rely on a priori assumptions on the nature of the lensing object (isolated galaxy, cluster member, etc.)—it requires the ability to unambiguously extract lensing features like well-defined multiple images, their shape and orientation, or individual clumps within lensed arcs (e.g., Wagner et al. 2018). In galaxy-scale strong lenses such as the simulated one of Fig. 1, the observed images take the form of relatively smooth highly distorted arcs (i.e., rings), for which it is error-prone to attempt extracting local features of multiply imaged regions of the source without performing full lens modeling (Galan et al. 2024). Nevertheless, the local lensing formalism can be expanded to lensed arcs with globally smooth surface brightness, as developed in Birrer (2021). Such an approach is well-suited to make use of the multiple elongated arcs observed around massive galaxy clusters, as a way to locally correct the global mass model (e.g., Yang et al. 2020).
@@ 8.3 Combining methods to mitigate systematic biases
We quantify the scatter associated with the choice of modeling method in Table 2, which records the mean and scatter of relevant source light and lens mass quantities among all the models. Although dependent on the type of data considered in this work, these numbers provide an estimation of systematic errors that one can expect from modeling imaging data with different methods. The systematic errors we quote here can be interpreted as lower bounds to those of a real-case scenario, since we placed ourselves in an idealized setting – no contamination by the lens light and perfect knowledge of noise properties and mass model family – which removes a subset of the known sources of biases or degeneracies. Ideally, the addition of such complicating factors may broaden the posterior distributions from individual models, thus making them statistically compatible (underlying systematic biases would then be unnoticeable). However, more realistically, the scatter between models is likely to increase as a consequence of more complex models potentially subject to different sources of biases. Performing similar analyses as the one presented here to a wider variety of strong lensing data (different resolution, S/N, lensing configuration, nature of the source, etc.) will help understand the generalization of our results. For this purpose, the framework we have developed should help and encourage the multiplication of such analyses on both simulated and real data sets.
@@
As we show in Fig. 6 and Table 2, combining together the results from an ensemble of methods using uniform weights removes the observed biases. While individual methods display systematic errors in different parameters—and not necessarily always for the same parameters and in the same directions—it is reassuring that overall, we do not observe a significant residual offset after combination. Only for the mass density slope ($\gamma$) one can observe a global trend towards larger values. We quantify in Table 2 the bias reduction between the biases from individual models and the one from the combined model. Among the eight mass model parameters, we find that an average bias reduction of 5.4, which is a substantial improvement on inferences from individual models alone. This result is reassuring and shows that analyzing a given data set using independent modeling methods is an efficient way to mitigate systematic errors.
In this work, we conservatively assumed equal weights when combining individual methods. This is similar to the recent work of Wong et al. (2024), where the authors combined measurements of $H_{0}$ obtained from by modeling the same lensed quasar with two different modeling codes, assuming equal weights. Nevertheless, we note that other approaches exist. The least conservative approach would be to multiply individual posteriors together, which is equivalent to assume that all methods are independent from each other and do not share systematic errors. The six modeling methods are only partially independent, as some of the models were performed by the same modelers using the same modeling code (see Appendix B) albeit with different source reconstruction techniques. Additionally, we modeled a single imaging data realization of our simulated strong lens, which further introduces some degree of statistical dependence. The most Bayesian approach would be to combine individual models based on their Bayesian evidence. However, for reasons we detailed in Sect. 8.1, objectively ranking models obtained from methods based on fundamentally different assumptions is still an unresolved issue. Nonetheless, better quantifying how different strong lens modeling methods perform on identical or similar datasets with respect to their complexity will be important to design better statistical combination procedures.
@@
Beyond the accuracy improvement of a combined posterior distribution over the lens model parameters, the comparison of multiple models—and if possible, truly independent models—is extremely valuable for detecting unknown sources of systematics, updating our current modeling assumptions and techniques and testing additional models with different levels of complexity. Similarly, in source plane, comparing the morphology of different versions of the unlensed object (which is never directly observable) is necessarily valuable for straightening the confidence in a given feature, which in turn improves the robustness of its interpretation. Moreover, different models may predict different observables that may otherwise be overlooked, enabling the potential detection of anomalous lens systems (such as missing images, Ertl et al. 2024) or anticipate the detection of additional images unobserved in current observations (e.g., similar to the geometric confirmation of multiple images in cluster-scale lenses, Diego et al. 2023).
@@ 8.4 Extrapolating to time-delay applications
Figure 12: Fermat potential differences evaluated at image positions of a multiply point source located at the center of the source galaxy. Given that the modeled data (Fig. 1) does not contain the multiply imaged point source, this can be seen as an intermediate scenario between that of a lensed quasar (i.e., centered on its host) and a lensed SN (i.e., has faded away).
@@
We show in Fig. 12 the Fermat potential differences for three hypothetical pairs of lensed images. While the simulated data we analyze in this work (Fig. 1) mimics the case of a lensed SN that faded away, we assume for simplicity that its location in source plane coincides with its host galaxy (although one could also proceed similarly for any other source position that leads to at least two lensed images). We label the lensed images ABCD, order them by ascending Fermat potential (i.e., $\Phi(\boldsymbol{\theta}_{A})$ is the lowest) and consider the three independent pairs $ij\in\{{\rm AB,AC,AD}\}$. In addition to the posterior mass parameters uncertainties, we add in quadrature an additional uncertainty to mimic the limited astrometric precision (for details, see Birrer & Treu 2019), assuming a conservative precision of 10 mas in image plane. The shaded gray area in each panels of Fig. 12 around the true values isolates the typical contribution of the astrometric scatter term, which is in our case small compared to the uncertainties modeling the extended lensed source. As is expected from the constraints on mass profile parameters (Fig. 6), we observe a similar scatter around the true values among the models.
Interestingly, we do not observe a clear correlation between model biases on individual lens potential parameters and their resulting biases on Fermat potential differences. For example, the Sparsity+Wavelets model, which display slight biases on some source properties and lens potential parameters (e.g., the lens position), give Fermat potential posteriors well-centered on the expected values. While a robust generalization of these results require further work (e.g., on a larger sample of simulations), the general trend we previously found still remains: the combination of individual posteriors assuming equal weights results in posteriors that are free of biases. These combined distributions are shown with dash-dotted black lines in Fig. 12, and encapsulate well the Fermat potential difference values. This further gives motivation for time delay cosmography analyses to compare and combine lens models together, in particular those obtained using independent modeling methods (see e.g., Shajib et al. 2022; Kelly et al. 2023). In a follow-up paper, we shall explore the dependency of systematic lens modeling errors on the source plane position of the time-delay background object.
@@ 8.5 Application to real data sets
For the purpose of demonstration, we focus on simulated imaging data, but the framework and ideas we present is intended to be applied to real data. The COOLEST standard, which provides a common ground to express lens models from different origins, already supports real data and their corresponding models (see e.g., Galan et al. 2024). Moreover, extending the standard to other types of models (for both the lens and the source) is straightforward. One complication for applying the same framework presented here to real data arises from the human time and computation time required to apply multiple codes and methods to a given data set (see also Sect. B).
@@
Although COOLEST significantly reduces the burden to express a lens model in a format that can be readily compared to other models, the acquisition of combined posterior distributions over lens model parameters still requires that at least two distinct lens models are in hand. Here we model the same data using four software packages and six modeling methods, which is likely unrealistic in most real scenarios, especially given the large amount of lenses that are still to be modeled (e.g., in the archival HST data, see Garvin et al. 2022) and that will be discovered in the near future (e.g., Collett 2015). Nevertheless, it is reasonable to assume that it is feasible to apply two or three methods to the same data (not necessarily within the same analysis), since the amount of lens modeling experts naturally grow over time and inference methods are becoming less and less time consuming. Deep learning methods, which we do not explore in the present work, offer promising avenues to accelerate the overall procedure either by proposing preliminary models to be refined with classical methods, or providing additional posterior distributions for final combination, at a negligible cost (ignoring the training phases).
When modeling real data, the chosen mass model parametrization is only an approximation of the true mass distribution. Extending the present analysis to account for differences between the truth and the model is obviously conceivable. However, it is important to note that such an extension inherently requires key assumptions in order to properly define “how different” is the truth compared to the model. A concrete example is the work of Gomer et al. (2022, Sect. 3.2) who simulated a population of SLACS-like strong lenses following a composite (baryons + dark matter) mass model, for which they had to assume a series of additional assumptions (ellipticity, halo scale radius, etc.), which they subsequently modeled with a power-law profile radially modified with a mass sheet parameter. Cao et al. (2022, Sect. 2.4) simulated strong lenses with mass distribution based on stellar kinematics models and performed modeling using single power-law profiles with shear. Van de Vyvere et al. (2022) quantified the biases caused by a lack of freedom along the azimuthal direction, when the true mass distribution contains multipoles or ellipticity gradients. The Rung 3 of the TDLMC, although focused on the recovery of $H_{0}$ instead of mass model parameters, has been a blind experiment where independent teams modeled lenses extracted from cosmological simulations (Ding et al. 2021). On real and simulated systems, the Time-Delay Cosmography collaboration (Millon et al. 2020) systematically uses both power-law and composite mass model families and marginalizes over the resulting posteriors, which are further constrained along the radial direction with stellar kinematics information (e.g., Shajib et al. 2022). Given the diversity of deviations to the commonly assumed sheared power-law elliptical mass profile in real galaxies—such as truncations (e.g., Limousin et al. 2007, in dense environments), cores (e.g., Collett et al. 2017) or multipoles (e.g., Stacey et al. 2024)—the generalization of the above-cited results to all strong lens systems is still unclear. It is possible that for a given real system, a set of dedicated simulations based on a set of likely mass models, followed by lens modeling as for the real data, may be the best way to properly quantify systematic errors caused by the choices of mass model.
@@ 9 Conclusion
We have conducted a fully blind modeling experiment on strong lensing data simulated with a dedicated software package by one author, while four other authors used four independent software packages to model it. In total, six modeling methods—which differ in their source reconstruction techniques and inference pipelines—have been applied on that same data. We have made a series of simplifying assumptions to keep this novel kind of analysis tractable, in particular regarding the lens light, the mass model parametrization and the noise properties. In contrast, the true PSF, lens parameters and source morphology were hidden and the optimization and inference strategies left free to the modelers. We used the lensing standard COOLEST to overcome the challenges that arise when comparing results obtained with different modeling codes. The resulting image and source plane models, as well as model residuals are given in Fig. 2, and constraints on lens potential parameters are shown in Fig. 6. Below we summarize our main results:
@@
  • We investigated how certain model assumptions affect the recovery of lens potential parameters. In particular, we explore (1) how the cuspy nature of the lensed galaxy can lead to systematic errors if the source model is not flexible enough to capture large intensity variations, and (2) how the accuracy of the PSF (the true PSF being unavailable to the modelers) plays a role even for extended source modeling. We find that both an accurately sampled PSF and a source model with large dynamic range (e.g., using a luminosity-weighted source prior) are warranted to reconstruct cuspy lensed sources while minimizing systematic errors on lens potential parameters.
Over the past years, numerous lens modeling methods have been proposed and implemented in different software packages. Here, we selected a subset of those with the goal of using them together, instead of only opposing them. Typically, we refrain from explicitly ranking the modeling methods, which would only be meaningful over an extremely large sample of strong lenses with different data quality and modeling assumptions to ensure proper generalization. In a real-case scenario, we do not have access to the truth; therefore, combining the results from different methods is a pragmatic and efficient way to detect and mitigate systematic errors. As is shown, for example, in Suyu et al. (2006) by testing three types of source regularizations, there exists an inherent dependence on the data properties, the lens configuration and the (unobservable) intrinsic source morphology, such that it is unlikely that a single source reconstruction method gives unbiased results in all cases. Our work strengthens this idea and goes further by combining a large collection of models, while investigating their specific effects on inferred parameters. Moreover, we purposefully allowed some level of freedom for the modelers (e.g., masks, PSF, posterior marginalization, etc.), such that our work also illustrates the role of specific modelers’ choices. These choices play a role in the observed scatter between models, and can be marginalized over by combining the methods together.